# H-STAR: LLM-driven Hybrid SQL-Text Adaptive Reasoning on Tables

### v1

@@ 3.2 Efficiency Analysis
H-STAR Column Retrieval: 2-4 6-10
Row Retrieval: 2-4
@@
4 Ablation Study
We conduct an ablation study to measure the contribution of each step in the pipeline and assess the impact of integrating symbolic-textual and adaptive reasoning in H-STAR.
4.1 Importance of Each Stage
To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study. This study involved systematically removing one stage at a time. For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process, while maintaining the table extraction stage. Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically by removing the row extraction and column extraction steps individually and then collectively. A detailed assessment of the contribution of each stage to the overall performance of the method is shown in Table 4. It highlights the importance of each step in the pipeline. Removing any step can result in a performance loss except for a marginal increase in the case of row extraction for TabFact.
@@ 4.1 Importance of Each Stage
H-STAR 85.92 68.51
w/o row extraction 86.17 66.30
w/o column extraction 84.04 67.03
w/o table extraction 83.79 63.58
w/o adaptive reasoning 79.35 61.47
Table 4: Performance on TabFact and WikiTQ datasets.
4.2 Impact of a Hybrid SQL-Text Approach
We perform a quantitative evaluation of the impact of the hybrid approach in H-STAR. Figures 5, 6 depict how using only one of the approaches results in insufficient outputs. Furthermore, it is evident from Table 5 that not employing a hybrid approach (H-STAR) results in a significant performance loss.
Method TabFact WikiTQ
CoT 79.05 60.43
H-STAR ${}_{\text{text}}$ 79.43 61.47
Text-to-SQL 52.05 42.03
H-STAR ${}_{\text{sql}}$ 58.50 46.09
H-STAR 85.93 68.51
The low accuracy of SQL approaches can be attributed to the incorrect final answers. The text-based reasoning approach achieves reasonably high scores on TabFact and WikiTQ, albeit lower than the hybrid approach. By combining textual and structured reasoning capabilities, the hybrid approach i.e., H-STAR can effectively handle diverse data formats and task requirements, leading to superior performance across various benchmarks.
5 Where do LLMs fail?
This section examines errors in H-STAR, contrasts the failure points with existing approaches, and summarizes our key findings.
5.1 Error Analysis on H-STAR
The disjoint, step-wise nature of H-STAR enables identifying and analyzing failures. Table 6 shows 100 randomly selected incorrect answer instances from the Tabfact and WikiTQ datasets. In this study, ‘Missing Columns’ and ‘Missing Rows’ refer to missing necessary columns and rows, respectively. ‘Incorrect Reasoning’ occurs when H-STAR extracts the correct table, but LLM fails to produce the correct answer. ‘Incorrect Annotations’ include semantically identical answers in different formats, ambiguous questions, and incorrect gold answers. Table 6 shows that for 100 TabFact failures, 2% are missing columns, 9% missing rows, 79% incorrect LLM reasoning, and 10% incorrect annotations. For WikiTQ, 6% are missing columns, 17% missing rows, 54% incorrect reasoning, and 23% incorrect annotations.
Tabfact WikiTQ
Missing Columns 2% 6%
Missing Rows 9% 17%
Incorrect Reasoning 79% 54%
Incorrect Annotation 10% 23%
Table 6: Error analysis on 100 error samples across datasets showcasing the distribution of errors for H-STAR.
Analysis. Fewer errors in column and row extraction indicate that H-STAR efficiently extracts the necessary table. By reducing errors in the table extraction process, we effectively "shift left" addressing problems earlier in the pipeline and enhancing overall performance.
5.2 More Error Analysis: H-STAR vs Others
Table 7 compares errors in H-STAR, with TabSQLify, and BINDER. We use 100 WikiTQ samples where TabSQLify fails. Many errors with TabSQLify (6/100 missing columns, 56/94 missing rows) result from incorrect row extraction. Out of 100, TabSQLify correctly extracts the table for 38 samples, but incorrectly reasons on 29 (76%). BINDER does not extract tables but generates multiple neural SQL queries for table reasoning tasks.
TabSQLify BINDER H-STAR
Missing Columns 6/100 N/A 4/100
Missing Rows 56/94 N/A 16/96
Incorrect Reasoning 29/38 70/100 27/80
Incorrect Annotation 9/9 9/30 10/53
Correct Answers 0/0 21/21 43/43
Table 7: Analysis of error types in 100 samples from WikiTQ where TabSQLify fails. Note: The denominator refers to the remaining samples; we start with 100 samples and subtract the number of error samples to determine the remaining samples.
However, its output shows that it fails to reason correctly in 70 out of 100 cases. In contrast, H-STAR proves more effective in both table extraction and reasoning, achieving correct generations in 43 samples. Additionally, it demonstrates fewer errors in table extraction (4/100 columns; 16/96 rows) and reasoning (27/80, i.e., 34% of samples).
5.3 What did we learn?
Firstly, our evaluation demonstrates that our hybrid approach achieves substantial improvements, surpassing the performance of previous state-of-the-art methods across various table reasoning tasks. Secondly, the quantitative analysis demonstrates that our ‘multi-view’ approach extracts tables with reduced irrelevant data, resulting in more concise tables. Furthermore, the qualitative analysis highlights fewer errors in our table extraction method compared to prior approaches, confirming a decrease in extraneous information.
Thirdly, our analysis indicates consistent performance even with longer tables, emphasizing our method’s effectiveness in accurately extracting relevant information by filtering out noise. The adaptive reasoning component boosts performance, especially for longer tables. Lastly, the ablation study shows that breaking down the task into sub-tasks significantly enhances the overall performance, with each sub-task playing a crucial role in achieving superior results. Moreover, it highlights the constraints of relying solely on either text or SQL approaches, which are effectively addressed by our H-STAR approach. Together, these findings emphasize the substantial advantages of our hybrid and multi-view approach in addressing complex table reasoning tasks.
6 Related Work
Decomposing problems into smaller, manageable tasks has proven effective in solving complex reasoning challenges Zhou et al. (2022); Khot et al. (2022). Recent techniques in table reasoning follow this approach, either by breaking tasks into fixed sub-tasks Cheng et al. (2022); Ye et al. (2023); Nahid and Rafiei (2024) or by employing iterative methods Jiang et al. (2023); Wang et al. (2023). BINDER Cheng et al. (2022) is an SQL-based approach that modifies SQL statements to include LLM API calls within SQL statements. DATER Ye et al. (2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning respectively. Chain-of-Table Wang et al. (2023) uses a textual reasoning approach to update tables iteratively. The final table post-update is used for final reasoning. Unlike H-STAR, these techniques depend either on symbolic or semantic reasoning but not both.
7 Conclusion
Limitations
Ethics Statement
Acknowledgments
Appendix A Additional Results
Tables 8 and 9 show results of other baseline models on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
Model Method Accuracy
PLM TaPas 48.8
GraPPa 52.7
LEVER 62.9
ITR 63.4
TableCoT 48.8
DATER 65.9
BINDER 61.9
ReAcTable 65.8
GPT-3.5-Turbo ReAcTable 52.5
TableCoT 52.4
StructGPT 52.2
GPT-3.5-Turbo H-STAR 68.8
Table 8: Comparison of H-STAR with additional methods on WikiTQ dataset.
Model Method Accuracy
PLM Table-BERT 68.1
LogicFactChecker 74.3
SAT 75.5
TaPas 83.9
TAPEX 85.9
SaMoE 86.7
PASTA 90.8
Codex TableCoT 72.6
DATER 85.6
BINDER 85.1
ReAcTable 83.1
GPT-3.5-Turbo ReAcTable 73.1
TableCoT 73.1
GPT-3.5-Turbo H-STAR 83.7
Prompting ROUGE-1 ROUGE-2 ROUGE-L
PaLM-2*
DATER 0.63 0.41 0.53
Chain-of-Table 0.66 0.44 0.56
GPT-3.5-Turbo
Table-CoT 0.62 0.39 0.51
Tabsqlify 0.58 0.35 0.48
H-STAR 0.62 0.39 0.52
Table 10: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model
Method Fluent Correct Adequate Faithful
T5-large 94.6 54.8 50.4 50.4
Human 95 92.4 95.6 95.6
TableCoT 96 82 75 87
Tabsqlify 97 88 84 93
H-STAR 96.6 87.6 89.6 94
Table 11: Human evaluation results on FeTaQA dataset.
A.1 FeTaQA
Human Evaluation
Qualitative Analysis
In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024). The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality. The answers are evaluated based on their fluency and adequacy when compared with the gold answers. Table 12 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify. ‘Both’ indicates samples where both the algorithms perform on par. This can be attributed to the use of the same LLM. Most of the failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction. Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table. It can be noted that H-STAR generates an answer comparable with the ground truth for the same. Figure 9 visually depicts the outputs from both methods. It illustrates the evaluation process undertaken to compare the quality of the outputs.
Method % of samples
Both 56%
TabSQLify 18%
H-STAR 26%
Table 12: Comparison of H-STAR with TabSQLify on 100 samples.
Figure 9: Comparison of the outputs with TabSQLify.
@@ B.2 Textual Reasoning
Appendix C Hyperparameters
Table 13 provides the details of PaLM-2 hyper-parameters used for the WikiTQ and TabFact datasets whereas Table 14 showcases the hyperparameter setting for GPT-3.5-Turbo. The hyperparameters ‘samples’ indicate the number of outputs taken for each step of the pipeline whereas the ‘examples’ indicate the number of few-shot demonstrations for each step.
@@ Appendix C Hyperparameters
Reasonsql 0.1 1.0 512 1 3
Reasontext 0.1 1.0 256 1 5
Table 13: Hyperparameter settings for PaLM-2
@@
Reasonsql 0.1 1.0 512 1 4
Reasontext 0.1 1.0 256 1 5
Table 14: Hyperparameter settings for GPT-3.5-Turbo

### v2

@@ A.3 Analysis: Generated Samples
H-STAR Column Extraction: 2-4 6-10
Row Extraction: 2-4
@@
A.4 Benefit of using Transposed Table
Table 14 compares H-STAR’s performance using transposed versus original tables for text-based column verification. On TabFact, H-STAR achieves 86.51% accuracy with transposed tables, versus 85.22% with original tables. On WikiTQ, accuracy is 68.62% with transposed tables, compared to 66.59% with original tables.
@@ A.4 Benefit of using Transposed Table
H-STAR (with transposed table) 86.51 68.62
H-STAR (with original table) 85.22 66.59
Table 14: Comparison between H-STAR using a) transposed table and b) original table for text-based column extraction.
Analysis. Transposing tables leads to better performance than using the original table for column verification. The transposed table approach is more effective for complex datasets like WikiTQ, which are harder to reason upon and contain longer tables.
@@ B.2 Textual Reasoning
Appendix C Comparison with ReAct
While our method appears to be similar to ReAct in terms of combining textual reasoning and tool-based methods, there are significant differences between H-STAR and ReAct:
1. Integration of Symbolic and Semantic Reasoning.
H-STAR combines SQL-based and text-based reasoning, while ReAct merges textual reasoning with actions. ReAct uses tools mainly for actions, not reasoning. H-STAR applies SQL for quantitative tasks and text for lexical queries, whereas ReAct focuses more on textual reasoning.
2. Fixed Multi-Stage Approach.
ReAct prompts LLMs to alternate between reasoning and actions, allowing interaction with external sources. H-STAR, on the other hand, splits tabular reasoning into two stages: 1) Table Extraction, and 2) Adaptive Reasoning, blending SQL and text reasoning. This hybrid approach sets H-STAR apart from ReAct’s focus on textual reasoning and interaction with the environment.
3. Table Extraction Approach.
H-STAR uses a "multi-view" approach for table extraction, identifying relevant columns from both original and transposed tables before selecting rows, reducing irrelevant data, and improving focus. In contrast, ReAct does not employ any extraction method.
4. Adaptive Reasoning Strategy.
H-STAR uses SQL to support semantic reasoning for quantitative tasks while relying solely on semantic reasoning for lexical queries. In contrast, ReAct interleaves reasoning and actions without adapting to task type. H-STAR’s hybrid approach, combining SQL logic and text understanding, excels in handling diverse tasks, outperforming ReAct on table reasoning tasks.
Appendix D Implementation Details
We provide the implementation details including the hyperparameters and prompt details for each of the individual steps. H-STAR operates in two key stages: 1) Table Extraction and 2) Adaptive Reasoning. The Table Extraction phase includes column extraction (SQL-based: colsql and text-based: coltext) and row extraction (SQL-based: rowsql and text-based: rowtext). The Adaptive Reasoning phase consists of SQL-based (fsql) and text-based (ftext) reasoning components.
D.1 Hyperparameters
Table 15 provides the details of PaLM-2 hyper-parameters used for the WikiTQ and TabFact datasets. Table 16 showcases the hyperparameter setting for GPT-3.5-Turbo. The hyperparameters ‘samples’ indicate the number of outputs taken for each step of the pipeline whereas the ‘examples’ indicate the number of few-shot demonstrations for each step.
@@ D.1 Hyperparameters
fsql 0.1 1.0 512 1 3
ftext 0.0 1.0 512 1 5
Table 15: Hyperparameter settings for PaLM-2.
@@
fsql 0.1 1.0 512 1 4
ftext 0.0 1.0 512 1 5
Table 16: Hyperparameter settings for GPT-3.5-Turbo.
D.2 Input Prompts
Figures 10, 11, 12, 13, 14 and 15 illustrate the input prompts containing sample in-context learning examples for the steps colsql, coltext, rowsql, rowtext, fsql, and ftext respectively.
Figure 10: Prompt for SQL-based column extraction (colsql).
Figure 11: Prompt for Column-based column verification (coltext).
Figure 12: Prompt for SQL-based row selection (rowsql).
Figure 13: Prompt for text-based row verification (rowtext).
Figure 14: Prompt for SQL-based reasoning (fsql).
Figure 15: Prompt for text-based reasoning (ftext).
Appendix E Error Analysis: Case Study
Figures 16 and 17 showcase how semantic and symbolic approaches can individually fail for the column and row extraction steps. Figures 19, 18, 20 and 21 illustrate error cases for each individual step within H-STAR.
Figure 18: Case Study: ‘Missing Columns’. An illustration of the error ‘Missing Columns’. In the above example, the extracted table does not contain the column ‘Tournament’ that contains the ‘Year End Ranking’. The omission of the ‘Tournament’ column affects the downstream steps in the pipeline leading to an insufficient row extraction and reasoning.
Figure 19: Case Study: ‘Missing Rows’. An illustration of the error ‘Missing Rows’. From the above example, it can be seen that while the extracted columns are correct, the row extraction is incorrect. Thus, as the extracted table lacks the necessary information, the final answer is incorrect.
Figure 20: Case Study: ‘Incorrect Reasoning’. An illustration of the error ‘Incorrect Reasoning’. Despite the correct question-relevant table being extracted, the model misinterprets the question. It adds up the scores by the opposing team instead thus returning an incorrect answer.
Figure 21: Case Study: ‘Incorrect Annotation’. An illustration of the error ‘Incorrect Annotation’. In the example, the query-specific table is correctly extracted and the final reasoning leads to the correct answer. However, the prediction is penalized for not being an ‘exact match’.